I have a problem that reads:
If f(x)=x^2 and g(x)=x-3, find (f (open circle) g)(0)
Could you please help me figure this out?

Answer by stanbon(75887)

I have a problem that reads:
If f(x)=x^2 and g(x)=x-3, find (f (open circle) g)(0)
Could you please help me figure this out?
------------------------------------------------
(fog)(x) = f[g(x)] = f[x-3] = (x-3)^2
Therefore (fog)(0) = (0-3)^2 = 9
======================================
Cheers,
Stan H.